OpenStudy (anonymous):

What is the solution of the system? Use substitution. x = –2y + 10 –6y = 6x – 6 z = –3x + 10y will award medal if helped

OpenStudy (anonymous):

hmmm. Did I go wrong somewhere?

OpenStudy (anonymous):

did u

OpenStudy (anonymous):

found it.

OpenStudy (anonymous):

Y = 9, sorry!

OpenStudy (anonymous):

x = -8

OpenStudy (anonymous):

Yes, and x = -8

OpenStudy (anonymous):

z=114

OpenStudy (anonymous):

and yes, for z just plug in the values.

OpenStudy (anonymous):

z = –3x + 10y x=-8, y = 9

OpenStudy (anonymous):

x = -2y+10 -6y = 6x-6 -6y = 6(-2y+10)-6 -6y=-12y+60-6 6y=54 y=9

OpenStudy (anonymous):

How many solutions does the system have? x = -4y + 4 2x + 8y = 8 another medal for this answer

OpenStudy (callisto):

consider the first and the second equation first x = –2y + 10 -(1) –6y = 6x – 6 -(2) Put (1) into (2) -6y = 6(–2y + 10) -6 -6y = -12y + 54 6y = 54 y=9 x= -2(9) +10 = -18 +10 = -8 put x=-8 and y=9 into z = –3x + 10y z= -3(-8) + 10(9) = 24+90 = 114

OpenStudy (anonymous):

It has 0 solutions

OpenStudy (anonymous):

im gonna repost another question the same one and answer it on that question so i can award the medal cool?

OpenStudy (callisto):

x = -4y + 4 -(1) 2x + 8y = 8 -(2) when (1)X2, we will get (2) --> not solvable since the 2 given equations are the same

OpenStudy (callisto):

or we should say infinitely many solutions

OpenStudy (anonymous):

its either no solutions or infinately many which one

OpenStudy (anonymous):

graphs of the functions coincide, so it has infinitely many solutions

OpenStudy (anonymous):

It's unsolvable... but you could say infinitely many solutions.

OpenStudy (anonymous):

anyone knows What is the solution of the system? Use elimination. 3x – 4y = 9 –3x + 2y = 9

OpenStudy (anonymous):

x = -9

OpenStudy (anonymous):

y = -9 . is this correct? :)

OpenStudy (anonymous):

yup